# Inverse Design in Nanophotonics via Representation Learning

### v1

@@ 3 Class B: Representation-Learning for the design space (Input-Side
The second class of inverse-design methods shifts the focus of ML from the solver or its output to the input or the geometry of the device. We interchangeably use the words input side and design space to denote the feasible space of the PDE-constrained optimization problem. Input-side representation learning seeks a latent manifold that captures the key low-dimensional features of feasible, high-performance structures. Rather than emulating the solver, these methods build a learnable representation of geometry. The objective is to discover low-dimensional manifolds (latent design spaces) that capture the important features of high-performance photonic structures [79, 80, 81, 82]. Central to this approach are techniques such as variational autoencoders (VAEs) [81], other dimensionality reduction and manifold learning methods [49], and generative adversarial networks (GANs) [52, 82, 83]. These models are typically trained on datasets comprising existing device geometries, paired with their corresponding optical responses.
Through this training, the models learn a compressed, structured representation, a latent space, where each point ideally corresponds to a unique, fabrication-consistent device geometry, as depicted in the (right-hand) Input-side branch of Figure 1. A generative model maps a small set of latent coordinates $\mathbf{z}$ to full two-dimensional (2D) or 3D layouts, making optimization more tractable while embedding practical design constraints. This cuts the dimensionality and the simulation cost but limits the exploration to shapes that the generator can express, a trade-off between efficiency and coverage. As a result, the learned latent manifold becomes a compact domain for subsequent design exploration and optimization.
@@
Generative models [85] are particularly prominent for design proposal and exploration within learned latent spaces. Note that in previous sections, although we put the emphasis on the wave-Y net model, GLOnet in Figure 2(e, left) is an input-side generative model [52]. Figures 3(a,b) depict a conditional adversarial autoencoder (c-AAE) pipeline for designing nanopatterned antennas [31]. In this system, the antenna topology, along with key geometric parameters like unit-cell size and spacer thickness, is encoded by an encoder network (E) into a compact 17-dimensional latent space (Figure 3(a)). A generator network (G) subsequently learns to produce realistic device patterns from vectors sampled from this latent space. A discriminator (D) keeps samples plausible and aligned with a chosen prior, yielding a physics-informed manifold. After training, the generator G is coupled with a pre-trained optical surrogate model (in this case, a visual geometry group (VGG)-based CNN predicting efficiency, as shown in Figure 3(b)) to enable rapid offline evaluation of proposed designs. The c-AAE’s primary function here is to learn an efficient representation of the input device designs—the latent space. New designs are generated by sampling or manipulating points within this learned input manifold.
@@
Another example, shown in Figures 3(c-e), employs a CNN-assisted WGAN for the inverse design of reconfigurable photonic waveguide arrays [84]. The device is a three-channel silicon rib-waveguide coated with the phase-change material antimony selenide (Sb2Se3); laser writing changes the material state and the coupling between channels (Figure 3c). A WGAN generator is trained to produce candidate phase-change patterns (representing the device geometry, Figure 3(e), left panel) from latent vectors $z_{i}$. Incidentally, a NN-based simulator then rapidly predicts the resulting transmission matrix of the device (an example field distribution is shown in Figure 3(d)). The differentiability of this simulator allows for the computation of performance gradients with respect to the latent variables $z_{i}$, guiding their iterative refinement to meet a specific design target (Figure 3(e), right panel). In this workflow, the WGAN learns to generate valid and potentially high-performing input patterns (the phase-change material configurations). Optimization is performed efficiently within the WGAN’s latent space, which is a learned, compressed representation of the design space.
@@
Hybrid approaches that combine learned input representations with standard optimizers are becoming common. The HiLAB framework (Figures 3(f,g)) integrates a VAE with Bayesian optimization (BO) for metagrating design [67]. A Vision Transformer-based VAE is first trained to encode 256×128 binary metagrating patterns into a low 8-dimensional latent space (Figure 3(f)). This process establishes a smooth, potentially fabrication-constrained, manifold for subsequent optimization. A Bayesian optimizer then operates in this learned geometric latent space concurrently with optimizing a small set of continuous physical hyperparameters, such as layer thicknesses ($t_{1}$,$t_{2}$) and the lattice constant ($\Lambda_{y}$) (Figure 3(g)). Each point suggested by the Bayesian optimizer in this joint space is decoded by the VAE’s decoder into a candidate metagrating design, which is then evaluated using full-wave EM simulations. Including training costs, this method enabled at least a ten-fold reduction in the number of simulations needed compared to a conventional TO, while achieving better performance to advance state-of-the-art. The subsequent optimization leverages this compressed latent space to efficiently search for designs that maximize a worst-case diffraction efficiency metric across multiple target wavelengths.
@@
The creation of a latent space through input-side learning serves as a powerful learned prior over the design space. This prior implicitly encodes information about good or physically plausible geometries. For instance, the c-AAE in Figure 3(a) [31] compresses designs into a 1-dimensional compressed design space, while the VAE in the HiLAB framework (Figure 3(f)[67]) establishes an 8-dimensional latent space described as a smooth, fabrication-constrained domain. Reducing the dimensionality may improve local optima. Optimization then proceeds by manipulating variables within these structured latent spaces, as seen in Figure 3(e) [84] and Figure 3(g) [67]. This prior guides the search, enhancing efficiency and increasing the likelihood of identifying valid, high-performing designs compared to unconstrained exploration in the full, high-dimensional parameter space. The structural properties of the latent space, such as smoothness or disentanglement of features, thus become critical factors for successful design outcomes.
@@
Additionally, input-side learning can facilitate knowledge discovery. The latent variables in a trained generative model can sometimes reveal latent design degrees of freedom [37] or correlate with specific physical properties or performance metrics, offering insights into underlying design principles that are not initially obvious.
Despite these strengths, input-side methods face hurdles. The first hurdle is the acquisition and creation of a sufficiently large, diverse, and high-quality training set of device designs and their associated performance metrics. Naive random sampling of the vast design space often results in a training dataset dominated by poor-performing structures. This leads to inefficient use of computational resources. This chicken-and-egg problem—whereby learning a good representation of high-performing designs ideally requires a dataset of such designs, which are often the very objects of the search—is a fundamental issue. To address this data scarcity, recent work uses hybrid data-generation schemes for training data generation. For instance, initial datasets might be seeded with designs found via early-stopped TO [67] or through exploration with low-fidelity surrogate models [88, 89], thereby focusing the representation learning process on more promising regions of the design space. This highlights an inherent interconnectedness and potential for synergy between different classes of inverse design methods. Further challenges include the representational capacity and potential biases introduced by the chosen model architecture (VAE, GAN, etc.). The learned manifold might not encompass all possible optimal designs, or it might be biased towards features prevalent in the training data. Finally, while some latent dimensions may correlate with interpretable physical features, achieving a full and intuitive understanding of all learned latent variables and their relationships is not guaranteed.
@@ 4 Global Search Methods Accelerated by Representation Learning
Global optimization schemes like GAs [64, 90], BO [91, 92, 93], and RL [57, 56, 55] aim to escape local optima by sampling or evolving designs across the feasible space. Unlike gradient-based design methods, which often converge to a local solution dependent on initial conditions, global methods systematically explore regions that gradient-based methods often miss. However, the main practical constraint is the computational budget required by these global optimization methods. Although theoretically capable of reaching global optima given infinite evaluation samples, methods such as GAs and RL quickly become computationally infeasible due to the necessity of full-wave EM simulations at each design evaluation. This computational cost scales exponentially [94] with the dimensionality of the design space, limiting the number of iterations and thus the exploration depth within realistic research scenarios.
@@
(a) Short-L3 InP nanobeam cavity with symmetric taper and mirror holes $(x_{1},x_{2},\ldots,x_{m})$. (b) User-specified optical targets are fed to the L2DO engine. (c) Deep-RL loop: a four-layer MLP policy (PPO/DQN) interacts with an FDTD environment; replay-buffered experience tuples $(s_{t},a_{t},r_{t},s_{t+1})$ guide optimization of hole positions, radii, and counts.
(d–f) Physics-informed RL (PIRL) for metagrating optimization. (d) Binary state encoding of a Si/SiO2(Si: silicon, SiO2: silicon oxide) meta-grating; objective—maximize first-order TM deflection efficiency $\eta$. (e) Physics-informed pre-training: a U-Net learns sensitivity maps $\Delta\eta_{\mathrm{approx}}$ from adjoint analysis (plot compares exact, adjoint-approximate, and NN-predicted $\Delta\eta$). (f) Parallel Deep-Q stage: the pretrained agent $Q_{0}^{\omega}$ is cloned into 16 workers running full-wave simulations; trajectories populate a global replay buffer while the master network $Q^{\omega}$ is synchronously updated, enabling efficient, large-scale exploration of the metagrating design space. Panels (a–c) reproduced/modified with permission from [55]; panels (d–f) reproduced/modified with permission from [56].
@@ 5 Concluding Remarks
One concern is how to handle and calibrate design complexity. Freeform topologies offer a large configuration space that can, in principle, yield superior performance at the risk of high computational costs and increased susceptibility to manufacturing errors [40]. More constrained parameterizations, such as shape-based [42] or level-set methods [102], reduce the search space. However, they potentially sacrifice some bandwidth or multi-resonant behavior. Future research may develop rigorous metrics or “complexity bounds” that quantify when adding more degrees of freedom ceases to provide meaningful gains, enabling more systematic choices about the granularity of optimization with the simplest parameterization that does not sacrifice performance.
Data availability is still an issue. Although hardware acceleration and modern frameworks have enabled more simulations, large and coherent public datasets for nanophotonics remain insufficient [103]. This gap contrasts with mainstream AI (artificial intelligence) fields and hinders purely ML-based design. Broad and shared databases, annotated with device geometries and their optical responses, could jump-start cross-laboratory studies, accelerate the development of advanced surrogate models, and facilitate systematic transfer learning. Without such resources, individual research groups must repeatedly generate data on a project-by-project basis, limiting reproducibility and collaboration. The dataset bootstrapping loop, wherein limited physics-based optimization generates initial data for an ML model, which then guides further design explorations, feeding new high-quality designs back to enhance the model, is useful for costly nanophotonic problems.
@@
A design representation may also embed desirable mathematical properties directly into learned representations to yield better-behaved optimization landscapes. Autoencoders, for example, cut the effective dimension, which often improves local minima. Furthermore, neural representations provide continuous relaxations for inherently discrete or binary design problems, such as TO, so gradient descent can still be leveraged even when the final design must be binary.
@@
A practical challenge is making sure that optimized structures are fabricable. Future work should bake fabrication constraints into the learned representation itself. Instead of constraining the minimum feature sizes or connectivity during the optimization, generative models could inherently produce fabricable designs directly. This approach draws inspiration from parameterizations, such as pillar-based designs with predefined size constraints. The critical choice is between implicitly teaching fabricability through data or explicitly enforcing it via parameterization or optimization objectives. Embedding fabrication rules in the latent space keeps the search inside fabricable shapes and cuts post-design rule checks.
Finally, many future devices will couple several physics, so multiphysics representations are needed. Current inverse design typically addresses a single physical domain (e.g., EMs), but advanced devices often involve coupled phenomena such as opto-mechanical, thermo-optic, or electro-optic effects. Unified representations that encode and optimize across coupled physics would enable truly multifunctional, robust devices. Approaches might include learning joint latent spaces influencing multiple physics or employing PINNs to solve coupled partial differential equations(PDEs). Multi-modal generative models could also play a vital role. Advancing beyond single-physics optimization to tackle coupled multiphysics represents a major research frontier.
@@
Progress on the numerical front is also reshaping inverse design. Faster integral-equation solvers [105, 106], GPU-accelerated PDE codes [107, 108, 109], and symmetry-aware solvers [110, 111] are utting simulation time. These faster solvers can mesh with iterative or hybrid algorithms and enlarge the tractable region of the search space, even if an exhaustive brute-force sweep remains out of reach. Frameworks that embed physical laws, such as Maxwell’s equations, directly into ML architectures as a solver layer may become practical to keep model outputs physically valid.
@@
In practice, selecting the best hybrid strategy for a given problem remains non-trivial. Common metrics for evaluating performance, such as objective function values or peak efficiencies, do not always capture robustness, fabrication constraints, or interpretability. Moreover, systematic guidelines for combining different modules (e.g., deciding when to switch from global search to local refinement) are still missing, highlighting the need for more research.
Looking ahead, it is clear that hybrid inverse design approaches—combining global exploration, local refinement, advanced solver strategies, and data-driven modeling—will continue to impact the development of nanophotonic devices. By integrating these elements, designers can balance between geometric freedom and manufacturing constraints, between computational feasibility and multi-functional objectives, and between empirical performance and fundamental physical limits. The end result may be a deeper, more unified understanding of how light and matter can be co-engineered in the nanoscale regime.

### v2

@@ 3 Class B: Representation-Learning for the design space (Input-Side
The second class of inverse-design methods shifts the focus of ML from the solver or its output to the input or the geometry of the device. We interchangeably use the words input side and design space to denote the feasible space of the PDE-constrained optimization problem. Input-side representation learning seeks a latent manifold that captures the key low-dimensional features of feasible, high-performance structures. Rather than emulating the solver, these methods build a learnable representation of geometry. The objective is to discover low-dimensional manifolds (latent design spaces) that capture the important features of high-performance photonic structures [83, 84, 85, 86].
@@
Central to this approach are techniques such as variational autoencoders (VAEs) [85], other dimensionality reduction and manifold learning methods [54], and generative adversarial networks (GANs) [57, 86, 88]. These models are typically trained on datasets comprising existing device geometries, paired with their corresponding optical responses.
Through training, the model learns a compact latent space in which each point corresponds to a unique, ideally fabrication-consistent geometry, as shown in the input-side branch of Figure 1. A generative model then maps a low-dimensional latent vector $\mathbf{z}$ to the full device layout. Far from being restrictive, this learned manifold serves as a powerful prior that regularizes the design search, particularly valuable in multi-objective problems where complete training data are unavailable [89]. A latent space trained on a subset of objectives can still guide optimization toward global solutions while avoiding out-of-distribution regions where surrogate predictions become unreliable. In this way, the input-side representation constrains the search to physically meaningful regions, complementing output-side learning in hybrid frameworks.
@@
Generative models [90] are particularly prominent for design proposal and exploration within learned latent spaces. Note that in previous sections, although we put the emphasis on the WaveY-Net net model, GLOnet in Figure 2(e, left) is an input-side generative model [57]. Figures 3(a,b) depict a conditional adversarial autoencoder (c-AAE) pipeline for designing nanopatterned antennas [31]. In this system, the antenna topology, along with key geometric parameters like unit-cell size and spacer thickness, is encoded by an encoder network (E) into a compact 17-dimensional latent space (Figure 3(a)). A generator network (G) subsequently learns to produce realistic device patterns from vectors sampled from this latent space. A discriminator (D) keeps samples plausible and aligned with a chosen prior, yielding a physics-informed manifold. After training, the generator G is coupled with a pre-trained optical surrogate model (in this case, a visual geometry group (VGG)-based CNN predicting efficiency, as shown in Figure 3(b)) to enable rapid offline evaluation of proposed designs. The c-AAE’s primary function here is to learn an efficient representation of the input device designs—the latent space. New designs are generated by sampling or manipulating points within this learned input manifold.
@@
Another example, shown in Figures 3(c-e), employs a CNN-assisted Wasserstein GAN (WGAN) [91] for the inverse design of reconfigurable photonic waveguide arrays [87]. The device is a three-channel silicon rib-waveguide coated with the phase-change material antimony selenide (Sb2Se3); laser writing changes the material state and the coupling between channels (Figure 3c). A WGAN generator is trained to produce candidate phase-change patterns (representing the device geometry, Figure 3(e), left panel) from latent vectors $z_{i}$. Incidentally, a NN-based simulator then rapidly predicts the resulting transmission matrix of the device (an example field distribution is shown in Figure 3(d)). The differentiability of this simulator allows for the computation of performance gradients with respect to the latent variables $z_{i}$, guiding their iterative refinement to meet a specific design target (Figure 3(e), right panel). In this workflow, the WGAN learns to generate valid and potentially high-performing input patterns (the phase-change material configurations). Optimization is performed efficiently within the WGAN’s latent space, which is a learned, compressed representation of the design space.
@@
Hybrid approaches that combine learned input representations with standard optimizers are becoming common. The HiLAB framework (Figures 3(f,g)) integrates a VAE with Bayesian optimization (BO) for metagrating design [71]. A Vision Transformer-based VAE is first trained to encode 256×128 binary metagrating patterns into a low 8-dimensional latent space (Figure 3(f)). This process establishes a smooth, potentially fabrication-constrained, manifold for subsequent optimization. A Bayesian optimizer then operates in this learned geometric latent space concurrently with optimizing a small set of continuous physical hyperparameters, such as layer thicknesses ($t_{1}$$t_{2}$) and the lattice constant ($\Lambda_{y}$) (Figure 3(g)). Each point suggested by the Bayesian optimizer in this joint space is decoded by the VAE’s decoder into a candidate metagrating design, which is then evaluated using full-wave EM simulations. Including training costs, this method enabled at least a ten-fold reduction in the number of simulations needed compared to a conventional TO, while achieving better performance to advance state-of-the-art. The subsequent optimization leverages this compressed latent space to efficiently search for designs that maximize a worst-case diffraction efficiency metric across multiple target wavelengths.
@@
The creation of a latent space through input-side learning serves as a powerful learned prior over the design space. This prior implicitly encodes information about good or physically plausible geometries. For instance, the c-AAE in Figure 3(a) [31] compresses designs into a 17-dimensional compressed design space, while the VAE in the HiLAB framework (Figure 3(f)[71]) establishes an 8-dimensional latent space described as a smooth, fabrication-constrained domain. Reducing the dimensionality may improve local optima. Optimization then proceeds by manipulating variables within these structured latent spaces, as seen in Figure 3(e) [87] and Figure 3(g) [71]. This prior guides the search, enhancing efficiency and increasing the likelihood of identifying valid, high-performing designs compared to unconstrained exploration in the full, high-dimensional parameter space. The structural properties of the latent space, such as smoothness or disentanglement of features, thus become critical factors for successful design outcomes.
@@
Furthermore, input-side learning can be an enabler of knowledge discovery. While the latent variables in a trained generative model are not always directly interpretable, techniques from the broader ML community offer pathways to enhance their physical meaning. Disentanglement methods, for instance, aim to learn representations where each latent dimension corresponds to a single factor of variation in the data [94, 95]. Frameworks like the $\beta$-TCVAE (Total Correlation VAE) explicitly penalize correlation between latent variables to encourage such factorial representations [95]. Applying these techniques to nanophotonic design could yield latent spaces where specific variables control distinct physical features (e.g., resonance location, feature size). This transforms the generative model from a black box into an intuitive design tool, as has been demonstrated in related fields like materials discovery [96], and substantiates its potential for knowledge discovery.
Despite these strengths, input-side methods face hurdles. The first hurdle is the acquisition and creation of a sufficiently large, diverse, and high-quality training set of device designs and their associated performance metrics. Naive random sampling of the vast design space often results in a training dataset dominated by poor-performing structures. This leads to inefficient use of computational resources. This chicken-and-egg problem—whereby learning a good representation of high-performing designs ideally requires a dataset of such designs, which are often the very objects of the search—is a fundamental issue. To address this data scarcity, recent work uses hybrid data-generation schemes for training data generation. For instance, initial datasets might be seeded with designs found via early-stopped TO [71] or through exploration with low-fidelity surrogate models [97, 98], thereby focusing the representation learning process on more promising regions of the design space. This highlights an inherent interconnectedness and potential for synergy between different classes of inverse design methods. Further challenges include the representational capacity and potential biases introduced by the chosen model architecture (VAE, GAN, etc.).
@@ 4 Global Search Methods Accelerated by Representation Learning
Practical inverse-design problems are often under-constrained or evolve dynamically (e.g., intelligent metasurfaces with shifting objectives or fabrication limits) [99]. Representation learning mitigates these challenges through a powerful, two-pronged approach: (i) input-side models provide a reusable, physics-agnostic manifold of device geometries, enabling efficient exploration and generation of new design candidates when constraints change (Section 3), while (ii) output-side models provide a fast surrogate solver, enabling rapid evaluation and re-optimization of those candidates against the new targets (Section 2). This transferability not only facilitates efficient adaptation to evolving requirements but also supplies the foundation for making subsequent global searches computationally tractable under strict simulation budgets. Leveraging these transferable representations, global optimization schemes such as GAs [68, 100], BO [101, 102, 103], and RL [62, 61, 60] aim to escape local optima by sampling or evolving designs across the feasible space. Unlike gradient-based design methods, which often converge to a local solution dependent on initial conditions, global methods systematically explore regions that gradient-based methods often miss. However, the main practical constraint is the computational budget required by these global optimization methods. Although theoretically capable of reaching global optima given infinite evaluation samples, methods such as GAs and RL quickly become computationally infeasible due to the necessity of full-wave EM simulations at each design evaluation. This computational cost scales exponentially [104] with the dimensionality of the design space, limiting the number of iterations and thus the exploration depth within realistic research scenarios.
@@
(a) Short L3 InP nanobeam cavity with symmetric taper and mirror holes $(x_{1},x_{2},\ldots,x_{m})$. (b) User-specified optical targets are fed to the L2DO engine. (c) Deep-RL loop: a four-layer MLP policy (PPO/DQN) interacts with an FDTD environment; replay-buffered experience tuples $(s_{t},a_{t},r_{t},s_{t+1})$ guide optimization of hole positions, radii, and counts. (d–f) Physics-informed RL (PIRL) for metagrating optimization. (d) Binary state encoding of a Si/SiO2(Si: silicon, SiO2: silicon oxide) meta-grating; objective—maximize first-order TM deflection efficiency $\eta$. (e) Physics-informed pre-training: a U-Net learns sensitivity maps $\Delta\eta_{\mathrm{approx}}$ from adjoint analysis (plot compares exact, adjoint-approximate, and NN-predicted $\Delta\eta$). (f) Parallel Deep-Q stage: the pretrained agent $Q_{0}^{\omega}$ is cloned into 16 workers running full-wave simulations; trajectories populate a global replay buffer while the master network $Q^{\omega}$ is synchronously updated, enabling efficient, large-scale exploration of the metagrating design space. Permissions: Panels (a–c) Reproduced under the terms of the CC BY 4.0 license.[60] Copyright 2023, The Author(s). Panels (d–f) Reproduced under the terms of the CC BY 4.0 license.[61] Copyright 2024, The Author(s).
@@ 5 Concluding Remarks
One concern is how to handle and calibrate design complexity. Freeform topologies offer a large configuration space that can, in principle, yield superior performance at the risk of high computational costs and increased susceptibility to manufacturing errors [41]. More constrained parameterizations, such as shape-based [43] or level-set methods [112], reduce the search space. However, they potentially sacrifice some bandwidth or multi-resonant behavior. Future research may develop rigorous metrics or complexity bounds that quantify when adding more degrees of freedom ceases to provide meaningful gains, enabling more systematic choices about the granularity of optimization with the simplest parameterization that does not sacrifice performance.
Data availability remains a critical bottleneck. To mitigate the high cost of full-wave simulations, hybrid data generation strategies are essential. Multi-fidelity training, which leverages numerous cheap, low-fidelity simulations (e.g., from approximate solvers) to inform a model trained on a few high-fidelity examples, is a promising direction [113, 114]. Similarly, active learning can intelligently select the most informative designs to simulate next, drastically reducing the data required to train an effective surrogate model [86, 105]. These strategies, coupled with community efforts to standardize and share datasets like MetaNet [115], are crucial for progress.
This focus on data generation leads to a deeper methodological question regarding its use: the utility of a physics-informed loss when training on data already generated by a high-fidelity solver. While such a loss is the central element for frameworks like MaxwellNet [80] that learn directly from the PDE, its added value in purely data-driven contexts is more debatable. The effectiveness of a physics-informed loss depends on the dimensionality and composition of the design space. As highlighted in [37], two main types of degrees of freedom govern the problem: physical variables governed by Maxwell’s equations and geometrical parameters derived from a choice of parameterization. When the number of physical degrees of freedom is small and the training data already comply with the governing equations, adding an explicit physics term offers limited benefit. In contrast, when both physical and geometrical degrees of freedom are large and strongly coupled, where the curse of dimensionality becomes significant, the inclusion of a PDE-based constraint provides a useful inductive bias that promotes physically consistent learning and improved generalization [64].
@@
A practical challenge is ensuring optimized structures are manufacturable. Future work should move beyond post-design checks and bake fabrication constraints into the learned representation itself. For instance, generative models could be inspired by methods that combine electromagnetic and structural topology optimization to produce self-supported, 3D-printable designs [127], or by level-set methods that inherently produce smooth, fabricable boundaries [128, 45].
Finally, advancing beyond single-physics optimization to tackle coupled multiphysics represents a major research frontier. Future devices will require co-design across domains like electromagnetics, thermodynamics, and mechanics. Promising frameworks are emerging to tackle this challenge. For instance, advanced formulations of PINNs are being developed specifically to solve tightly coupled thermo-mechanical systems [129]. In parallel, material-informatics-based inverse design has been used to efficiently co-optimize metamaterials for distinct physical objectives, such as achieving selective performance in both the visible and infrared spectra [130]. While still an emerging area, these pioneering works demonstrate the high potential of multiphysics inverse design for creating truly multifunctional, robust devices.
Further challenges include the representational capacity of the chosen model architecture. This is a particularly critical consideration with the rise of multifunctional devices, which inherently involve multi-objective optimization where the goal is to map out a Pareto front of optimal trade-offs [131, 132]. A single design might excel in one objective (e.g., high efficiency at one wavelength) at the expense of another, while a different design might offer balanced performance across all objectives; both are valid solutions along the front. Since generating a training set that uniformly covers this entire front is often computationally prohibitive, the data is frequently biased toward specific, high-interest regions. Therefore, the coverage of the learned manifold becomes a strategic element of the design process rather than a simple shortcoming [133, 134, 135]. This reality, however, raises critical questions for future work. For example, can transfer learning be used to leverage a model trained on one region of the Pareto front to efficiently explore another? Furthermore, it remains an open question whether a single or shared latent representation is sufficient to capture the complex topology of an entire front, or if more sophisticated, task-specific representations are necessary. Finally, a related challenge is that of interpretability; while some latent dimensions may correlate with intuitive physical features, achieving a comprehensive understanding of all learned variables and their relationships is not guaranteed. Addressing these questions is a key frontier for the field.
@@
Progress on the numerical front is also reshaping inverse design. Faster integral-equation solvers [136, 137], GPU-accelerated PDE codes [138, 139, 140], and symmetry-aware solvers [141, 142] are cutting simulation time. These faster solvers can mesh with iterative or hybrid algorithms and enlarge the tractable region of the search space, even if an exhaustive brute-force sweep remains out of reach. Frameworks that embed physical laws, such as Maxwell’s equations, directly into ML architectures as a solver layer may become practical to keep model outputs physically valid.
A central challenge in inverse design is the one-to-many problem, where multiple distinct designs yield the same output. While generative input-side models like cGANs can address this by producing a diverse family of plausible solutions, they are inherently limited to the scope of their training data. Hybrid frameworks provide powerful strategies to transcend this limitation and discover higher-performance or novel solutions through two key synergies. First, pairing an input-side model with a differentiable output-side surrogate enables gradient-based local refinement [87, 64, 143]. A design generated from the latent space serves as a high-quality starting point, which is then fine-tuned using gradients from the surrogate to converge to a superior local optimum. Second, coupling an input-side latent space with a global search algorithm is uniquely powerful for co-optimizing the device pattern alongside its physical parameters (e.g., layer thicknesses, periodicity) [71, 57, 144]. While a generative model can provide diverse patterns for a fixed engineering objective, a global optimizer on a low-dimensional latent space can efficiently explore the much larger joint space of geometry and physical dimensions. This is critical in photonics for discovering novel resonances that emerge from the interplay between pattern and structure, a task for which standalone generative models are ill suited  [145, 146]. This transforms the design process from a static selection of plausible patterns to a dynamic search for truly optimal, and often non-intuitive, devices.
@@
In practice, selecting the best hybrid strategy for a given problem remains non-trivial. Common metrics for evaluating performance, such as objective function values or peak efficiencies, do not always capture robustness, fabrication constraints, or interpretability. Moreover, systematic guidelines for combining different modules (e.g., deciding when to switch from global search to local refinement) are still missing, highlighting the need for more research. At the heart of this strategic challenge is the trade-off between the input-side prior and the output-side physics constraint. A highly regularized input-side prior can relax the accuracy requirements of the output-side surrogate, while a highly accurate surrogate enables a broader, less constrained search. Navigating the balance between these complementary forms of regularization is a fundamental consideration for the future of advanced inverse design.
Looking ahead, it is clear that hybrid inverse design approaches, uniting global exploration, local refinement, advanced solvers, and data-driven modeling, will continue to impact the development of nanophotonic devices. This integration allows designers to balance key trade-offs: geometric freedom versus manufacturing constraints, computational cost versus multi-functional objectives, and empirical performance versus fundamental physical limits. The ultimate expression of this synergy may be the emergence of autonomous multi-agentic frameworks that leverage these hybrid methods as a high-performance computational backend. Such frameworks can translate high-level, semantic design goals into optimized device layouts in a nearly real-time, automated fashion, fundamentally shifting the paradigm of scientific discovery [147]. Crucially, by enabling the optimization of specific near-field distributions and optical modes, these advanced frameworks move beyond simple performance maximization, providing deeper insights into the underlying physics of light-matter interactions [145]. The end result will be a more profound understanding of how to co-engineer light and matter at the nanoscale.